Find the force that acts on a particle with acts on a particle with a charge of 3.2 ×10-19C travelling at a velocity of 5.5 ×103m/s perpendicular to a magnetic field with a flux density of 60mT
1
Expert's answer
2020-04-30T10:37:31-0400

A charged particle moving in a magnetic field is affected by the Lorentz force

(1)"\vec F=q[\vec V х\vec B]" In our problem are given "q=3.2\cdot 10^{-19}C" , "V=|\vec V|=5.5\cdot10^3m/s" , "B=|\vec B|=60 mT=60\cdot 10^{-3}T=6.0\cdot 10^{-2}T" . The vector cross product in (1) can be written as

(2) "[\vec V х\vec B]=\hat n VBsin(\alpha)" , where "\hat n" is an unit vector with direction perpendicular to "\vec V" and "\vec B" defined by the right hand rule ("\hat k=[\hat iх\hat j])" ), "\alpha" is the angle between two vectors "\vec V" and "\vec B" . Since we are given that the particle travelling with a velocity perpendicular to a magnetic field "\alpha=90\degree" and "sin(\alpha)=1" . If we choose a coordinate system so that "\vec V=V\cdot \hat i" and "\vec B=B\cdot \hat j" Lorentz force will be

(3) "\vec F=\hat k \cdot qVB=\hat k \cdot 3.2\cdot 10^{-19}\cdot 5.5 \cdot 10^3\cdot 6.0\cdot 10^{-2}=\hat k \cdot 1.056\cdot 10^{-16}N"

Answer: The force that acts on a particle is "\simeq1.1\cdot 10^{-16}N"